# Funnel Synthesis via LMI Copositivity Conditions for Nonlinear Systems

### v2

@@ 2.3 Invariance condition
for the system (12) with the feedback gain $K(t)\in\mathbb{R}^{n_{u}\times n_{x}}$. With this linear feedback, $\xi$ becomes linear in $\eta$. It follows that $\eta\in\mathcal{E}_{Q}$ implies $\xi\in\mathcal{E}_{KQK^{\top}}$ [17] where
@@
We refer $\mathcal{E}_{KQK^{\top}}$ to as an input funnel. Notice that if $KQK^{\top}$ is PD, the set $\mathcal{E}_{KQK^{\top}}$ is equivalent to $\{\xi\mid\xi^{\top}(KQK^{\top})^{-1}\xi\leq 1\}$. Using the state and input funnels, a funnel $\mathcal{F}(t)$ is defined as an Cartesian product of state and input funnel centered around the nominal state $\bar{x}(t)$ and input $\bar{u}(t)$, respectively, that can be written as
@@
for almost all $t\in[t_{0},t_{f}]$. Then, with every almost everywhere continuous signal $w(\cdot)$ such that (2), $\mathcal{E}_{Q}(t)$ defined in (16) is an invariant time-varying ellipsoid for the system (12) on $[t_{0},t_{f}]$, that is, if $\eta(\cdot)$ is a solution of (12) with $\eta(t_{0})\in\mathcal{E}_{Q}(t_{0})$, then $\eta(t)\in\mathcal{E}_{Q}(t)$ for all $t\in[t_{0},t_{f}]$.
@@
While Lemma 1 follows a similar proof as in [19, Lemma 1], it extends the result to consider more general types of disturbance signals $w(\cdot)$ that are continuous almost everywhere. In the next lemma, we define the invariance of the funnel $\mathcal{F}$ and its relevance to the invariance of $\mathcal{E}_{Q}$.
@@
If $\mathcal{E}_{Q}(t)$ is invariant for the uncertain LPV system (12) with the linear feedback control $\xi=K\eta$, then the funnel $\mathcal{F}(t)$ defined in (19) is invariant for the original system (1) in a sense that if $x(\cdot)$ is a solution of (1) with $x(t_{0})\in\{\bar{x}(t)\}\oplus\mathcal{E}_{Q}(t)$ and the control law,
@@ 2.4 Differential linear matrix inequality
It follows that
@@
We omit the time arguments in Lemma 3 and its proof for the notational brevity. The matrices $C,\,D,\,E,\,G$, and the constant $\lambda_{w}$ are time-invariant, while the terms $\tilde{A},\,\tilde{B},\,\tilde{F},\,Q,\,Y,\,K,\,W,\,N_{1}$, and $N_{2}$ are time-varying. Notice that with the fixed positive constant $\lambda_{w}$, $D_{LMI}(\dot{Q},Q,Y,\lambda_{\beta},\lambda_{\gamma})$ in (26) is linear in its arguments.
@@ 3.1 Transformation of DLMI into a finite number of LMIs
To transform the DLMI (26) into a finite number of LMIs, we first parameterize $Q(t)$, $Y(t)$, $\lambda_{\beta}(t)$, and $\lambda_{\gamma}(t)$ by emplyoing the same FOH interpolation used in (8), that is,
~$\tilde{\square}(t)$ $\coloneqq\sigma_{1}^{k}(t)\square_{k}+\sigma_{2}^{k}(t)\square_{k+1},\quad t\in[t_{k},t_{k+1}],$
@@
for all $k$ in $\mathbb{Z}_{[0,N-1]}$. These constants $(\gamma_{i})_{k}$ and $\beta_{k}$ serve as upper bounds for $\gamma_{i}(t)$ and $\beta(t)$, respectively, on each subinterval $[t_{k},t_{k+1}]$. Hence, they become valid constants for (7) and (10), respectively on each subinterval $[t_{k},t_{k+1}]$. This allows us to incorporate $(\gamma_{i})_{k}$ and $\beta_{k}$ as constants in the LMIs associated with each subinterval, facilitating the derivation of a finite set of LMIs for the entire horizon $[t_{0},t_{f}]$
@@
The similar results were discussed in [21, 3]. Here, we can further derive a less conservative LMI condition implying $P\in\mathcal{C}$ compared to (40).
@@
The matrix-valued function $Q(t)$ with the FOH (8) is not differentiable at each temporal node points $t=t_{k}$ for all $k$ in $\mathbb{Z}_{[0,N]}$. This does not compromise our objective of continuous-time invariance since it suffices for the Lyapunov condition (20) to hold for almost all $t$ in $[t_{0},t_{f}]$ according to Lemma 1.
@@ 3.2 Constraints
$\mathcal{P}_{x}$ =$=\{x\mid(a_{i}^{h})^{\top}x\leq b_{i}^{h},\quad i=1,\ldots,m_{x}\},$ (42a)
$\mathcal{P}_{u}$ =$=\{x\mid(a_{j}^{g})^{\top}u\leq b_{j}^{g},\quad j=1,\ldots,m_{u}\}.$ (42b)
@@
00 0 $\preceq\left[\begin{array}[]{cc}\left(b_{i}^{h}-(a_{i}^{h})^{\top}\bar{x}_{k}\right)^{2}&(a_{i}^{h})^{\top}Q_{k}\\ Q_{k}a_{i}^{h}&Q_{k}\end{array}\right],i=\mathbb{Z}_{[1,m_{x}]},$ (43c)
00 0 $\preceq\left[\begin{array}[]{cc}\left(b_{j}^{g}-(a_{j}^{g})^{\top}\bar{u}_{k}\right)^{2}&(a_{j}^{g})^{\top}Y_{k}\\ Y_{k}^{\top}a_{j}^{g}&Q_{k}\end{array}\right],j=\mathbb{Z}_{[1,m_{u}]}.$ (43f)
@@ 3.3 Problem formulation
The feasibility condition in (43) enforces the constraints only at the temporal node points, which could result in violations between these node points. A possible remedy is to increase the number of node points. It is important to note that while the feasibility is only enforced at each node point, the invariance holds across the entire horizon $[t_{0},t_{f}]$ as established in Theorem 1.
@@ 3.4 Estimation of the constants via sampling
To solve the problem (44), it is necessary to obtain the local Lipschitz constant $(\gamma_{i})_{k}$ and the bound $\beta_{k}$ described in (10) for each $k\in\mathbb{Z}_{[0,N-1]}$. If the system is globally Lipschitz, one can directly use the global Lipschitz constant for all $k$ [22]. One effective way for estimating local $\gamma_{k}$ is to use sampling approaches [3], collecting samples of the triple $(\eta_{s},\xi_{s},w_{s})$ around the nominal trajectory, with $s$ denoting each sample. Specifically, this involves sampling the state deviation $\eta_{s}$ within the maximum state funnel $Q_{max}$, the input deviation $\xi_{s}$ within the constraint (42b), and the disturbance within its bound (2) for each subinterval $[t_{k},t_{k+1}]$. To obtain a less conservative constant, one might use the iterative procedure provided in [3] where the estimation of $\gamma_{k}$ and funnel computation are alternately repeated until the convergence. This paper does not consider this iterative process for simplicity as our proposed method can be incorporated with any way to estimate the Lipschitz constant. On the other hand, the value of $\Delta(t)$ defined in (6) is known and can thus be evaluated at each $t$. Hence, for each subinterval $[t_{k},t_{k+1}]$, we pick $N_{s}$ sample temporal points $t_{s}$ uniformly, and then $\beta_{k}$ can be determined through
$\beta_{k}=\max_{s=1,\ldots,N_{s}}\lVert\Delta(t_{s})\rVert_{2}.$
@@ 4 Numerical simulation
We consider the unicycle model given in (4) with $c_{1}=0.03$ and $c_{2}=0.05$. We divide a $15$ second time horizon into $N=20$ subintervals, starting at $t_{0}=0$ and ending at $t_{f}=15$. The cost function $J$ in (44) is set as $-\text{trace}(Q_{0})+\text{trace}(Q_{N})$ to maximize the funnel entry and minimize the funnel exit. The nominal trajectory illustrated in Fig. 2 starts at $(0,0)$, passes through $(4,8)$, and ends at $(8,0)$. A state constraint for avoiding an obstacle is taken into account, resulting in a nonconvex state constraint as shown in Fig. 2. This nonconvex constraint is linearized around the nominal trajectory to generate the polytopic constraints in the form of (42). The constraints for inputs are defined as follows: $0\leq u_{1}\leq 2$ (m/s) and $\lvert u_{2}\rvert\leq 1.5$ (rad/s). The constant $\lambda_{w}$ is chosen as $0.3$ and the parameter $Q_{max}$ for the maximum funnel size is set to $\text{diag}(2^{2},2^{2},(20\pi/180)^{2})$. To estimate $(\gamma_{i}){k}$ and $\beta_{k}$, we sample 100 triples of $(\eta_{s},\xi_{s},w_{s})$ and 20 samples of $\Delta(t_{s})$ for each subinterval $[t_{k},t_{k+1}]$. The Mosek solver is used to solve the SDP (44). The code for the simulation can be found in the following repository.
@@
The results of the state funnel and input funnel are provided in Fig. 2 and Fig. 3, respectively. It is clear that state and input constraints are satisfied at each node point. We conduct a comparison between two different conditions, (40) and (41), in terms of the resulting cost and computation time for synthesizing the funnel. The funnel computed using (41) has smaller cost, at -1.670, compared to -1.615 for the funnel using (40). The computational time for MOSEK to solve the SDP (44) with (40) is longer, at 8.34s, compared to 1.64s with (41), which is aligned with our expectation.
@@
Finally, to test the invariant condition of the funnel, we take a total of 100 samples from the surface of the funnel entry $Q_{0}$. Then, these samples are propagated through the control law (21). During the propagation, we randomly choose the disturbance $w$ such that $\lVert w\rVert_{2}=1$. In the bottom figure of Fig. 4, the value of the Lyapunov function $V$ for each sample over the time horizon is illustrated. We can see that the values maintain below one, which means all trajectories remain in the funnel by the invariance property.

### v3

@@ 2.3 Invariance condition
for the system (12) with the feedback gain $K(t)\in\mathbb{R}^{n_{u}\times n_{x}}$. With this linear feedback, $\xi$ becomes linear in $\eta$. It follows that $\eta\in\mathcal{E}_{Q}$ implies $\xi\in\mathcal{E}_{KQK^{\top}}$ where
@@
We refer $\mathcal{E}_{KQK^{\top}}$ to as an input funnel. We use the subscript $KQK^{\top}$ because, when $KQK^{\top}$ is PD, the set $\mathcal{E}_{KQK^{\top}}$ is equivalent to $\{\xi\mid\xi^{\top}(KQK^{\top})^{-1}\xi\leq 1\}$. When $KQK^{\top}$ is only PSD (not PD), $\mathcal{E}_{KQK^{\top}}$ becomes a degenerated ellipsoid, which remains well-defined and compatible with (18). Using the state and input funnels, a funnel $\mathcal{F}(t)$ is defined as an Cartesian product of state and input funnel centered around the nominal state $\bar{x}(t)$ and input $\bar{u}(t)$, respectively, that can be written as
@@
for almost all $t\in[t_{0},t_{f}]$. Then, with every almost everywhere continuous signal $w(\cdot)$ such that (2), $\mathcal{E}_{Q}(t)$ defined in (16) is an invariant time-varying ellipsoid for the system (12) on $t_{0},t_{f}]$, that is, if $\eta(\cdot)$ is a solution of (12) with $\eta(t_{0})\in\mathcal{E}_{Q}(t_{0})$, then $\eta(t)\in\mathcal{E}_{Q}(t)$ for all $t\in[t_{0},t_{f}]$.
@@
For a detailed discussion of similar Lyapunov conditions in the context of input-to-state stability, we refer the reader to [23]. The similar Lyapunov condition with the time-invariant Lyapunov function is also studied in [24, Lemma 1]. While Lemma 1 follows a similar proof, we extend the result to consider more general types of disturbance signals $w(\cdot)$ that are continuous almost everywhere. In the next lemma, we define the invariance of the funnel $\mathcal{F}$ and its relevance to the invariance of $\mathcal{E}_{Q}$.
@@
If $\mathcal{E}_{Q}(t)$ is invariant for the uncertain LPV system (12) with the linear feedback control $\xi=K\eta$, then the funnel $\mathcal{F}(t)$ defined in (19) is invariant for the original system (1) in a sense that if $x(\cdot)$ is a solution of (1) with $x(t_{0})\in\{\bar{x}(t_{0})\}\oplus\mathcal{E}_{Q}(t_{0})$ and the control law,
@@ 2.4 Differential linear matrix inequality
By pre- and post-multiplying by $\eta^{\top},p^{\top},w^{\top}]$ and its transpose, respectively, and applying the identity $\dot{Q}^{-1}=-Q^{-1}\dot{Q}Q^{-1}$, we obtain
@@
We omit the time arguments in Lemma 3 and its proof for the notational brevity. The matrices C,D,E,GC,\,D,\,E,\,G, and the constant $\lambda_{w}$ are time-invariant, while the terms $\tilde{A},\,\tilde{B},\,\tilde{F},\,Q,\,Y,\,K,\,W,\,N_{1}$, and $N_{2}$ are time-varying. Notice that with the fixed positive constant $\lambda_{w}$, $D_{LMI}(\dot{Q},Q,Y,\lambda_{\beta},\lambda_{\gamma})$ in (26) is linear in its arguments.
@@ 3.1 Transformation of DLMI into a finite number of LMIs
To transform the DLMI (26) into a finite number of LMIs, we first parameterize $Q(t)$, $Y(t)$, $\lambda_{\beta}(t)$, and $\lambda_{\gamma}(t)$ by employing the same FOH interpolation used in (8), that is,
$\tilde{\square}(t)$ $\coloneqq\sigma_{1}^{k}(t)\square_{k}+\sigma_{2}^{k}(t)\square_{k+1},\quad t\in[t_{k},t_{k+1}],$
@@
for all $k$ in $\mathbb{Z}_{[0,N-1]}$. These constants $(\gamma_{i})_{k}$ and $\beta_{k}$ serve as upper bounds for $\gamma_{i}(t)$ and $\beta(t)$, respectively, on each subinterval $t_{k},t_{k+1}]$. Hence, they become valid constants for (7) and (10), respectively on each subinterval $t_{k},t_{k+1}]$. This allows us to incorporate $(\gamma_{i})_{k}$ and $\beta_{k}$ as constants in the LMIs associated with each subinterval, facilitating the derivation of a finite set of LMIs for the entire horizon $t_{0},t_{f}]$
@@
Similar results were discussed in [26, 4]. Here, we can further derive a less conservative LMI condition implying $P\in\mathcal{C}$ compared to (40).
@@
The matrix-valued function $Q(t)$ with the FOH (8) is not differentiable at each temporal node points $t=t_{k}$ for all $k$ in $\mathbb{Z}_{[0,N]}$. This does not compromise our objective of continuous-time invariance since it suffices for the Lyapunov condition (20) to hold for almost all $t$ in $t_{0},t_{f}]$ according to Lemma 1.
@@ 3.2 Constraints
$\mathcal{P}_{x}$ $=\{x\mid(a_{i}^{h})^{\top}x\leq b_{i}^{h},\quad i=1,\ldots,m_{x}\},$ (42a)
$\mathcal{P}_{u}$ $=\{x\mid(a_{j}^{g})^{\top}u\leq b_{j}^{g},\quad j=1,\ldots,m_{u}\}.$ (42b)
@@
$0$ $\preceq\left[\begin{array}[]{cc}\left(b_{i}^{h}-(a_{i}^{h})^{\top}\bar{x}_{k}\right)^{2}&(a_{i}^{h})^{\top}Q_{k}\\ Q_{k}a_{i}^{h}&Q_{k}\end{array}\right],i=\mathbb{Z}_{[1,m_{x}]},$ (43c)
$0$ $\preceq\left[\begin{array}[]{cc}\left(b_{j}^{g}-(a_{j}^{g})^{\top}\bar{u}_{k}\right)^{2}&(a_{j}^{g})^{\top}Y_{k}\\ Y_{k}^{\top}a_{j}^{g}&Q_{k}\end{array}\right],j=\mathbb{Z}_{[1,m_{u}]}.$ (43f)
@@ 3.3 Problem formulation
The feasibility condition in (43) enforces the constraints only at the temporal node points, which could result in violations between these node points. A possible remedy is to increase the number of node points. It is important to note that while the feasibility is only enforced at each node point, the invariance holds across the entire horizon $t_{0},t_{f}]$ as established in Theorem 1.
Remark 3.
The feasibility of the problem (44) depends on many factors such as the magnitude of $\gamma$ and $\beta$, the density of the node points, the configuration and size of obstacles, and the bounds on control inputs. Due to this complexity, it is generally difficult to guarantee feasibility a priori. However, a key advantage of the proposed approach is that it is formulated as a convex optimization problem, certifying the problem’s infeasibility. In particular, if the problem is infeasible, the solver can provide a formal certificate of infeasibility.
@@ 3.4 Estimation of the constants via sampling
To solve the problem (44), it is necessary to obtain the local Lipschitz constant $(\gamma_{i})_{k}$ and the bound $\beta_{k}$ described in (10) for each $k\in\mathbb{Z}_{[0,N-1]}$. If the system is globally Lipschitz, one can directly use the global Lipschitz constant for all $k$ [27]. One effective way for estimating local $\gamma_{k}$ is to use sampling approaches [4], collecting samples of the triple $(\eta_{s},\xi_{s},w_{s})$ around the nominal trajectory, with $s$ denoting each sample. Specifically, this involves sampling the state deviation $\eta_{s}$ within the maximum state funnel $Q_{max}$, the input deviation $\xi_{s}$ within the constraint (42b), and the disturbance within its bound (2) for each subinterval $t_{k},t_{k+1}]$. To obtain a less conservative constant, one might use the iterative procedure provided in [4] where the estimation of $\gamma_{k}$ and funnel computation are alternately repeated until the convergence. This paper does not consider this iterative process for simplicity as our proposed method can be incorporated with any way to estimate the Lipschitz constant. On the other hand, the value of $\Delta_{e}(t)$ defined in (6) is known and can thus be evaluated at each $t$. Hence, for each subinterval $t_{k},t_{k+1}]$, we pick $N_{s}$ sample temporal points $t_{s}$ uniformly, and then $(\beta_{e})_{k}$ can be determined through
$(\beta_{e})_{k}=\max_{s=1,\ldots,N_{s}}\lVert\Delta_{e}(t_{s})\rVert_{2}.$
3.5 Algorithm summary
The overall procedure for synthesizing the funnel and controller is summarized below.
Algorithm 1 Funnel Synthesis
1:Input: Nominal trajectory, system dynamics, state/input constraints
2:Estimate $\gamma$ and $\beta$ as described in Section 3.4
3:for each $\lambda_{w}$ in the search grid do
4:  Solve the SDP problem in (44)
5:  Store the cost and feasibility status
6:end for
7:Select $\lambda_{w}$ with the lowest feasible cost
8:Return the corresponding funnel and controller
@@ 4 Numerical simulation
In this section, the proposed method is validated through two robotic motion planning and control applications: one for a unicycle model and the other for six-degree-of-freedom (6-DoF) powered descent guidance. For both examples, the Mosek solver is used to solve the SDP (44). The code for the simulation can be found in the following repository.
@@
4.1 Unicycle model
We consider the unicycle model given in (4) with $c_{1}=0.03$ and $c_{2}=0.05$. We divide a $15$ second time horizon into $N=20$ subintervals, starting at $t_{0}=0$ and ending at $t_{f}=15$. The cost function $J$ in (44) is set as $\text{trace}(Q_{0})+\text{trace}(Q_{N})$ to maximize the funnel entry and minimize the funnel exit. The nominal trajectory illustrated in Fig. 2 starts at $(0,0)$, passes through $(4,8)$, and ends at $(8,0)$. A state constraint for avoiding an obstacle is taken into account, resulting in a nonconvex state constraint as shown in Fig. 2. This nonconvex constraint is linearized around the nominal trajectory to generate the polytopic constraints in the form of (42). The constraints for inputs are defined as follows: $0\leq u_{1}\leq 2$ (m/s) and $\lvert u_{2}\rvert\leq 1.5$ (rad/s). The parameter $Q_{max}$ for the maximum funnel size is set to $\text{diag}(2^{2},2^{2},(20\pi/180)^{2})$. To estimate $(\gamma_{i}){k}$ and $\beta_{k}$, we sample 100 triples of $(\eta_{s},\xi_{s},w_{s})$ and 20 samples of $\Delta(t_{s})$ for each subinterval $t_{k},t_{k+1}]$.
@@ 4.1 Unicycle model
Figure 4: The cost results of (44) with different values of $\lambda_{w}$ are presented for both cases: using (40) in Lemma 4 and (41) in Lemma 5.
@@
Figure 6: (Top) Evolution of the Lyapunov function for the funnel computed using the baseline approach. (Bottom) Time history of the maximum eigenvalue of $D_{LMI}$, as defined in (26), for the funnel obtained by both the baseline and proposed methods.
@@
To determine a value for $\lambda_{w}$, we solved the SDP problem (44) across a range of candidate $\lambda_{w}$ values. The resulting cost values are summarized in Fig. 4, which shows that the minimum cost is attained when $\lambda_{w}=0.26$. The results of the state funnel and input funnel are provided in Fig. 2 and Fig. 3, respectively. It is clear that state and input constraints are satisfied at each node point. We conduct a comparison between two different conditions, (40) and (41), in terms of the resulting cost and computation time for synthesizing the funnel. The funnel computed using (41) has smaller cost, at -3.652, compared to -3.600 for the funnel using (40). The computational time for MOSEK to solve the SDP (44) with (40) is longer, at 11.63s, compared to 1.57s with (41), which is aligned with our expectation. Furthermore, Fig. 4 shows that for every case, the cost associated with (41) is consistently lower than that obtained by (40).
@@
To test the invariance condition of the funnel, we take a total of 500 samples from the surface of the funnel entry $Q_{0}$. Then, these samples are propagated through the control law (21). During the propagation, we randomly choose the disturbance $w$ such that $\lVert w\rVert_{2}=1$. In the bottom figure of Fig. 5, the value of the Lyapunov function $V$ for each sample over the time horizon is illustrated. We can see that the values maintain below one, which means all trajectories remain in the funnel by the invariance property.
We compare the proposed method with a baseline approach that directly discretizes the DLMI (26) using the forward Euler method. This type of discretization has been commonly used in related funnel computation literature [3, Section 4.2]. To validate the baseline approach, we follow the same procedure as with the proposed method: 500 samples are taken from the surface of the funnel entry computed using the baseline, and each sample is propagated under the control law. The evolution of the Lyapunov function for each sample is shown in the top of Fig. 6. Some trajectories exhibit Lyapunov function values greater than 1, indicating that the funnel computed by the baseline method violates the invariance condition. To explicitly assess satisfaction of (26) over time, we plot the time history of the maximum eigenvalue of $D_{LMI}$, which is defined in (26), using the decision variables from both the baseline and proposed methods in Fig. 6. The results clearly show that the baseline approach violates the DLMI condition between the node points, whereas the proposed method satisfies it throughout the entire time horizon.
4.2 6-DoF Powered Descent Guidance
We consider a 6-DoF rigid-body rocket model in an East-North-Up inertia coordinate described by
$\left[\begin{array}[]{c}\dot{p}\\ \dot{v}\\ \dot{\Phi}\\ \dot{\Omega}\end{array}\right]$ $=\left[\begin{array}[]{c}v\\ \frac{1}{m}C_{I/B}(\Phi)F+g\\ R(\Phi)\Omega\\ J^{-1}(T+r\times F-\Omega\times J\Omega)\end{array}\right],$
where the state vector $x=(p,v,\Phi,\Omega)$ consists of the position $p\in\mathbb{R}^{3}$, velocity $v\in\mathbb{R}^{3}$, Euler angles $\Phi\in\mathbb{R}^{3}$, and angular velocity $\Omega\in\mathbb{R}^{3}$, and the control input $u=(F,T)$ includes the thrust force $F\in\mathbb{R}^{3}$ and torque $T\in\mathbb{R}^{3}$. The constant $m\in\mathbb{R}$ is the vehicle mass, $J\in\mathbb{R}^{3\times 3}$ is the inertia matrix, and $g\in\mathbb{R}^{3}$ is the gravitational constant. The vector $r\in\mathbb{R}^{3}$ is the position vector from the center of the mass to the location of the thrust. The matrix $C_{I/B}(\Phi)$ is the rotation matrix from the body frame to the inertia frame and the matrix $R(\Phi)$ represents the transformation between Euler angle rates and body-frame angular velocity. More details in the derivation of the model could be found in [28].
The nominal trajectory has a plane maneuver in the plane defined by $r_{x}=r_{y}$, with $r_{x}$ and $r_{y}$ denoting the position components along the East and North axes, respectively. We use nondimensionalized units, with $U_{T}$, $U_{L}$, and $U_{M}$ denoting the units of time, length, and mass units, respectively. The mass $m$ is set to 22 $(U_{M})$ and the gravity is set to $g=[0,0,-1.625]^{\top}$ $(U_{M}/U_{T}^{2})$. We divide an entire horizon into $N=5$ subintervals, starting at $t_{0}=0$ $(U_{T})$ and ending at $t_{f}=7.457$ $(U_{T})$. The cost function is set to the same as in the unicycle example: $\text{trace}(Q_{0})+\text{trace}(Q_{N})$. The constraint sets $\mathcal{X}$ and $\mathcal{U}$ are defined as
$\mathcal{X}$ $=\{x\mid x_{lb}\leq x\leq x_{ub}\},$
$\mathcal{U}$ $=\{u\mid u_{lb}\leq u\leq u_{ub}\},$
$x_{ub}$ $=(8,8,8,2,2,2,30\degree,30\degree,30\degree,45\degree,45\degree,45\degree),$
$x_{lb}$ $=-x_{lb},$
$u_{ub}$ $=(1.5,1.5,5.0,0.1,0.1,0.1),$
$u_{lb}$ $=-(1.5,1.5,0.0,0.1,0.1,0.1).$
Since the bounded disturbance $w$ is not considered in this example, it is not necessary to perform the line search for $\lambda_{w}$; thus $\lambda_{w}$ is set to zero and the matrix $F(t)$ is set to the zero matrix for all time.
The computed funnels projected onto the two dimensional East-North plane and the three-dimensional East-North-Up space are illustrated in Fig. 7 and in Fig. 8, respectively. The funnel computed using (41) has a cost of -31.982, which is lower than the cost of the funnel computed using (40), -31.890. The MOSEK solver takes 28.10 seconds to solve the SDP with (41) and 17.41 seconds with (40).
Similar to the unicycle example, 500 samples are taken from the funnel entry, and each sample is propagated under the synthesized control law. Each sample’s trajectory in the position coordinates is illustrated in Fig. 8. From the results, we observe that although the nominal trajectory has an in-plane maneuver, the resulting funnel and sampled trajectories exhibit out-of-plane maneuver. The funnel invariance is validated by examining the evolution of the Lyapunov function, as shown in Fig. 10, confirming that all trajectories remain inside the funnel. The resulting input funnel, projected onto each input dimension, is given in Fig. 9, demonstrating that the input constraints are satisfied. The state funnel projected onto the Euler angle dimensions is given in Fig. 10. While the funnel satisfies the state constraints at each node point, slight constraint violations occur between nodes.
Figure 7: The synthesized state funnel projected onto the position coordinates. Filled ellipsoids represent the funnels at discrete node points, while the intermediate funnels between nodes are shown as unfilled ellipsoids.
Figure 8: (Top) The computed funnel projected onto the three-dimensional position space. The funnel is shown in blue, and the propagated sample trajectories are shown in purple. (Bottom) Evolution of the Lyapunov function of the sample trajectories.
Figure 9: The synthesized input funnel onto each input dimension.
Figure 10: The synthesized state funnel projected onto each Euler angle dimension.